Personal Electronix sells computer tablets and MP3 players. The business is divided into two divisions along product lines. CVP income statements for a recent quarter’s activity are presented below. Tablet Division MP3 Player Division Total Sales $854,000 $546,000 $1,400,000 Variable costs 631,960 300,300 932,260 Contribution margin $222,040 $245,700 467,740 Fixed costs 136,981 Net income $330,759 Determine sales mix percentage and contribution margin ratio for each division. Sales Mix Percentage Tablet division % MP3 Player division % Contribution Margin Ratio Tablet division % MP3 Player division % LINK TO TEXT Calculate the company’s weighted-average contribution margin ratio. (Round computations and final answer to 2 decimal places, e.g. 15.26%.) Weighted-average contribution margin ratio % LINK TO TEXT Calculate the company’s break-even point in dollars. (Round computations to 2 decimal places and final answer to 0 decimal places, e.g. 1,526.) Break-even point $ LINK TO TEXT Determine the sales level in dollars for each division at the break-even point. (Round computations to 2 decimal places and final answers to 0 decimal places, e.g. 1,526.) Break-even point Tablet division $ MP3 Player division $ Click if you would like to Show Work for this question: Open Show Work
Answer #1

Solution :

Computation of Sales mix and CM ratio
Particulars Tablet Division MP3 Player division Total
Sales $854,000.00 $546,000.00 $1,400,000.00
Variable costs $631,960.00 $300,300.00 $932,260.00
Contribution margin $222,040.00 $245,700.00 $467,740.00
Sales mix (%) 61.00% 39.00%
Contribution margin ratio 26.00% 45.00% 33.41%

Breakeven point in dollar sales = Fixed costs / Weighted average contribution margin ratio

= $136,981 / 33.41% = $410,000

Sales level for tablet division at break even point = $410,000 * 61% = $250,100

Sales level for MP3 Player division at break even point = $410,000 * 39% = $159,900